# A Comprehensive Survey on 3D Content Generation

### 02 Feb 2024 v1

@@ 3.3 Human Avatar
Previously, most 3D native generative methods were confined to constrained datasets like ShapeNet containing only fixed object categories. Recent advances in text-to-image diffusion models new possibilities. DreamFusion  Poole et al. (2023) leveraged score distillation sampling techniques to transfer knowledge from powerful 2D diffusion models into optimizing 3D representations like NeRF, achieving a significant boost in text-to-3D synthesis quality. This paradigm rapidly expanded the scope of diffusion-based approaches from objects to other domains such as scenes and humans. Several milestone methods are presented in Figure 2.
@@ 4.1 Object
DreamFusion pioneered the paradigm of optimizing a unique 3D representation per text input or per image, guided by powerful pretrained 2D diffusion models. This approach established a new foundation but also revealed key challenges ahead - namely, achieving high-fidelity quality in resolution, geometric detail and texture fidelity; ensuring consistent generation across diverse views, known as the ”multi-face Janus problem”; and optimizing synthesis speed for interactive applications.
@@ 4.2 Scene
The main idea of 2D prior-based scene generation is to leverage 2D pretrained large model to generate partial scenes. Then, inpaiting strategy is applied to generate large-scale scenes. Text2room Höllein et al. (2023) is a typical example of using 2D pretrained model to generate depth of 2D image. Then, the image is inpainted with more depths. These depths are merged to generate large scale scene. LucidDreamer Chung et al. (2023) first generates multi-view consistent images from inputs by using inpaiting strategy. Then, the inpainted images are lifted to 3D space with estimated depth maps and aggregate the new depth maps into the 3D scene. SceneTex Chen et al. (2023a) generates scene textures for indoor scenes using depth-to-image diffusion priors. The core of this method lies in the proposal of a multiresolution texture field that implicitly encodes the appearance of the mesh. The target texture is then optimized using VSD loss in respective RGB renderings. Additionally, SceneDreamer Chen et al. (2023c) introduces a Bird’s Eye View (BEV) scene representation and a neural volumetric renderer. This framework learns an unconditional generative model from 2D image collections. With this model, it becomes possible to generate unbounded 3D scenes from noise without any specific conditioning.
@@ 5.1 Object
While the newly generated views in Zero123 were consistent with the given view, consistency was not maintained between generated novel views. In response, several works aimed to produce multiple views simultaneously with inter-view consistency. SyncDreamer Liu et al. (2024), MVDream  Shi et al. (2024) all enabled generating multiple perspectives at once, with information exchange between views to ensure consistency. Wonder3D  Long et al. (2023) introduced a normal modal and fine-tuned a multi-view Stable Diffusion model to concurrently output RGB and normal maps across perspectives. One-2-3-45++  Liu et al. (2023a) advanced multi-view 3D generation via an enhanced Zero123 module enabling simultaneous cross-view attention, alongside a multi-view conditioned 3D diffusion module performing coarse-to-fine textured mesh prediction over time.
@@ 5.4 Dynamic
Recent works attempted to reconstruct 3D scenes based on generated videos, introducing a new 4D pipeline that generates a video and then complements its 3D representation. 4DGen Yin et al. (2023) made pseudo multi-view videos via multi-view diffusion prior and optimized the reconstruction of gaussian splattings based on multi-resolution hex-plane. DreamGaussian4d Ren et al. (2023) deployed 3D-aware diffusion prior to supervise the multi-view reconstruction of a given video and refined the corresponding scene with video diffusion prior.

### 19 Mar 2024 v2

@@ 3.3 Human Avatar
Previously, most 3D native generative methods were confined to constrained datasets like ShapeNet containing only fixed object categories. Recent advances in text-to-image diffusion models and gan offer new possibilities. Several milestone methods are presented in Figure 2.
@@ 4.1 Object
DreamFusion Poole et al. (2023) pioneered the paradigm of optimizing a unique 3D representation per text input or per image, guided by powerful pretrained 2D diffusion models. This approach established a new foundation but also revealed key challenges ahead - namely, achieving high-fidelity quality in resolution, geometric detail and texture fidelity; ensuring consistent generation across diverse views, known as the ”multi-face Janus problem”; and optimizing synthesis speed for interactive applications.
@@ 4.2 Scene
In most 2D prior-based scene generation approaches, the primary method is to utilize pretrained large models to generate partial scenes. Subsequently, an inpainting strategy is employed to generate large-scale scenes. Text2room Höllein et al. (2023) is a typical example of using 2D pretrained model to generate depth of 2D image. Then, the image is inpainted with more depths. These depths are merged to generate large scale scene. LucidDreamer Chung et al. (2023) first generates multi-view consistent images from inputs by using inpaiting strategy. Then, the inpainted images are lifted to 3D space with estimated depth maps and aggregate the new depth maps into the 3D scene. SceneTex Chen et al. (2023a) generates scene textures for indoor scenes using depth-to-image diffusion priors. The core of this method lies in the proposal of a multiresolution texture field that implicitly encodes the appearance of the mesh. The target texture is then optimized using VSD loss in respective RGB renderings. Additionally, SceneDreamer Chen et al. (2023c) introduces a Bird’s Eye View (BEV) scene representation and a neural volumetric renderer. This framework learns an unconditional generative model from 2D image collections. CityDreamer Xie et al. (2024)brings nature scene generation to the city level, enhancing the realism of city generation by decoupling buildings and background generation.
@@ 5.1 Object
While the newly generated views in Zero123 were consistent with the given view, consistency was not maintained between generated novel views. In response, several works aimed to produce multiple views simultaneously with inter-view consistency. SyncDreamer Liu et al. (2024), MVDream  Shi et al. (2024) all enabled generating multiple perspectives at once, with information exchange between views to ensure consistency. Wonder3D  Long et al. (2023) introduced a normal modal and fine-tuned a multi-view Stable Diffusion model to concurrently output RGB and normal maps across perspectives. One-2-3-45++  Liu et al. (2023a) advanced multi-view 3D generation via an enhanced Zero123 module enabling simultaneous cross-view attention, alongside a multi-view conditioned 3D diffusion module performing coarse-to-fine textured mesh prediction over time. UniDream Liu et al. (2023f) incorporates diffusion priors to generate multi-view albedo-normal information, which is further processed by a trained reconstruction model to produce 3D representations.
@@ 5.4 Dynamic
Recent works  Jiang et al. (2023) attempted to reconstruct 3D scenes based on generated videos, introducing a new 4D pipeline that generates a video and then complements its 3D representation. 4DGen Yin et al. (2023) made pseudo multi-view videos via multi-view diffusion prior and optimized the reconstruction of gaussian splattings based on multi-resolution hex-plane. DreamGaussian4d Ren et al. (2023) deployed 3D-aware diffusion prior to supervise the multi-view reconstruction of a given video and refined the corresponding scene with video diffusion prior.